OpenStudy (anonymous):

the sum of three consecutive integers is 51. find the product of the integers

OpenStudy (studygurl14):

Okay, so 3 consecutive integers. Do you know what that means?

OpenStudy (anonymous):

not a clue

OpenStudy (studygurl14):

3 consecutive integers means 3 numbers, one after another. Example: 2, 3, 4 Those are 3 consecutive integers. What's the differences between 2 and 3, and 2 and 4? If 2 = N, then 3 would be (N + 1) and 4 would be (N + 2), right? Use this information to create an equation for your problem

OpenStudy (texaschic101):

nice explaining StudyGurl :)

OpenStudy (anonymous):

i am still lost.....

OpenStudy (studygurl14):

Thanks @texaschic101

OpenStudy (anonymous):

Consecutive means oneright after another, right. So it's asking for three numbers, that are one after the other; let x be the first number, then the next number- x+1; x+2 so the three cinsecutive numbers would be; x, x+1, x+2 so we have; Find 3 consecutive integers with the sum of 51. x+x+1+x+2=51 add like terms; 3x+3=51 3x=48 x=16 x+1=17 x+2=18

OpenStudy (texaschic101):

so much for explaining it :(

OpenStudy (studygurl14):

Yep, what @Lemaz put. Except, they are asking for the product of the integers also

OpenStudy (anonymous):

but that is not it tho. it is not one of the options.

OpenStudy (anonymous):

the options are A. 4,080 B. 3,360 C. 4,020 D. 4,896

OpenStudy (texaschic101):

product means multiply...so multiply your 3 terms : 16 * 17 * 18 = your answer

OpenStudy (anonymous):

D